OpenStudy (anonymous):

Solve the equation for m: A = 2(m - w) thank u so much

OpenStudy (anonymous):

How about distributing the 2? So, \(\ \huge A=2m-2w \)

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

divide by 2: \[\large \frac {A}{2}=m-w \] then add w: \[\large \frac{A}{2}+w=m\]